# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ Abstract.
We consider the configuration space of points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system generates an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229. All authors are partially supported by JST CREST Grant No. JPMJCR1911.
@@ 1. Introduction
Given a simple finite graph $(V,E)$ with edge weights $\ell:E\to\mathbb{R}$, its $d$-dimensional realisation is a map $\phi:V\to\mathbb{R}^{d}$ that satisfies the condition $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$. This mapping embeds the graph into $d$-dimensional Euclidean space while preserving the prescribed edge lengths. The collection of all possible realisations forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$. This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [10, 21, 25]. The primary motivation for this work is rooted in the analysis of the motion of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. A state of a Kaleidocycle corresponds to a realisation of the graph formed by the vertices and edges of the tetrahedra (Eq. (2.1)), where $\ell$ is determined by the shape of the tetrahedron. Its motion, disregarding self-collision, is described by a path in the configuration space. Although this object has been studied by several authors [8, 28, 30], even its existence beyond simple classical cases has not been clarified. Our work addresses this gap by providing an explicit construction of the object and its motion.
@@
Our key observation is to associate each state of a Kaleidocycle with a closed spatial polygonal curve. This curve is formed by connecting the midpoints of the hinge edges. The angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the curvature angle (§2.1). In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle. We can then identify the configuration space of a given Kaleidocycle with a subspace of a direct product of two-spheres (2.2). Note that it is not clear if there exists $\ell$ (the shape of the tetrahedra) for a given $k$ such that the configuration space is non-empty; that is, if a Kaleidocycle exists. We construct a one-dimensional subspace of the configuration space using elliptic theta functions, proving that there is a Kaleidocycle for any $k\geq 6$ that can move. The idea is to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
@@
It was known in [19] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations [19] are defined on the configuration space (see §3.2). Therefore, a one-dimensional orbit in the configuration space is generated through the time evolution defined by the system of ordinary differential equations. The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [11]. We show that our construction of the motion of the Kaleidocycle (§5) indeed satisfies the system of the ordinary differential equations described in [19]. The sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 16].
@@
Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, Kaleidocycle. It provides an interesting example, where a flow generated by an integrable system on a real-algebraic set is explicitly integrated using elliptic theta functions.
@@
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. The state of a Kaleidocycle is identified with a solution to a system of quadratic equations corresponding to the geometric constraints on the binormal vectors. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. A condition for the resulting curve to be closed is also discussed; for a given $k$, the appropriate choice of parameters yields a Kaleidocycle. By introducing an additional parameter, a family of time-parametrised Kaleidocycles is constructed in Section 5. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the Kaleidocycles and show connections to previous research. Section 6 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present two conjectures that are suggested by numerical experiments. The Kaleidocycles constructed in this paper seem to be Möbius Kaleidocycles (Conjecture 6.1).
Most of the proofs of our theorems and propositions are deferred to the appendix since they require technically-involved computations regarding elliptic theta functions.
@@ 2. Algebraic and geometric models of Kaleidocycle
2.1. Configuration space of Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and makes an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [29] and the references therein).
We first define a generalisation of Kaleidocycle following [18]. We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring. Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces. Let $l_{01},l_{12},l_{20}$ be the lengths of the sides of the triangle face. A state of a $k$-Kaleidocycles of type $(l_{01},l_{12},l_{20})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k-1})\in(\mathbb{R}^{3})^{2k}$ of the quadratic equations
@@ 2.1. Configuration space of Kaleidocycle
$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12}.$
@@
Note that collision between tetrahedra is ignored. The solution space of (2.1) admits an action of the Euclidean transformation group $E(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $63\cdot 2k-5k-\dim(E(3))=k-$ 6. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [8]. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 00, giving an example of an overconstrained  [8]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [20].Some of the findings of [20] are also published in [30]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of underconstrained linkage mechanisms whose configuration space has less dimension than the expected one.
The state of a Kaleidocycle can be equivalently described by a spatial polygon or a closed discrete spatial curve. Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k-1)$ be the mid-point of the hinge edge. We set $\gamma_{k}=\gamma_{0}$. Since the tetrahedron is equifacial, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $\widetilde{B}_{n}$ and $\widetilde{B}_{n+1}$, where $\widetilde{B}_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$ is the hinge direction. By reordering the vertices, we can assume the inner product $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ is constant for $1\leq n<k$. Denote this constant by $\cos\lambda$ with $0<\lambda<\pi$, which we call the torsion angle. We have two cases $\langle\widetilde{B}_{k-1},\widetilde{B}_{0}\rangle=\pm\cos\lambda$. We say the Kaleidocycle is oriented (resp. non-oriented) when the value is $+\cos\lambda$ (resp. $-\cos\lambda$). We set $\widetilde{B}_{k}=\pm\widetilde{B}_{0}$ so that $\langle\widetilde{B}_{k-1},\widetilde{B}_{k}\rangle=\cos\lambda$. Furthermore, since $\widetilde{B}_{n+1}\times\widetilde{B}_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, we have $\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. Conversely, the vertex coordinates $x_{n}$ can be reconstructed from $\{\gamma_{n}\}$, $\{\widetilde{B}_{n}\}$ and $(l_{01},l_{12},l_{20})$ up to translation. Therefore,
Definition 1.
The configuration space of the $k$-Kaleidocycles with the torsion angle $\lambda$ is the set $\{(\widetilde{B}_{0},\ldots,\widetilde{B}_{k})\in(S^{2})^{k+1}\}$ satisfying the following conditions:
(2.2) $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ $=\cos\lambda,\quad(1\leq n\leq k)$
$\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})$ $=0,$
$\widetilde{B}_{k}$ =±$=\pm\widetilde{B}_{0}.$
A motion of a Kaleidocycle is identified with a one-dimensional (time-parametrised) orbit in the space defined by (2.1) or (2.2).
@@
Figure 1. (Left) The classical 6-Kaleidocycle. (Right) Kaleidocycle as a discrete curve with its framing.
@@ 2.2. Discrete spatial curve of a constant torsion angle
A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals. Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle\neq 0$ for all $n\in\mathbb{Z}$. A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows. Define the tangent $T:\mathbb{Z}\rightarrow S^{2}$ and the normal $\widetilde{N}:\mathbb{Z}\rightarrow S^{2}$ by
@@
(2.5) $\begin{split}\langle T_{n},T_{n-1}\rangle&=\cos\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
@@
Note that our Frenet frame coincides with the usual one only up to sign. We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes the state when the curvature vanishes.
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen. The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By Călugăreanu-Fuller-White’s theorem [1], the writhing number of the curve should stay constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.

### v3

# An explicit construction of Kaleidocycles by elliptic theta functions

@@ Abstract.
We consider the configuration space of ordered points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they simultaneously satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra greater than five. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system is explicitly solved to generate an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
Key words and phrases:
Kinematic chain, Configuration space, Discrete differential geometry, Integrable systems, Elliptic theta function
2000 Mathematics Subject Classification:
Primary 53A04, 53A70; Secondary 53A17, 70B15, 37K25, 37K10, 35Q53
The first-named author was partially supported by KAKENHI, Grant-in-Aid for Scientific Research (B) 25K00921 and (S) 25H00399. The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229.
@@ 1. Introduction
Given a natural number $k$ and a real number $1<\cos\lambda<1$, we consider the following space of configurations of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
(1.1) $\mathcal{C}^{\pm}_{k,\lambda}=\{(b_{0},\ldots,b_{k})\in(S^{2})^{k+1}\mid$ $\langle b_{n-1},b_{n}\rangle=\cos\lambda,\,(1\leq n\leq k),$
$\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=0,$
$b_{k}=\pm b_{0}\}.$
Our main result states that for any $k\geq 6$, there exists $\lambda$ such that a circle can be embedded in $\mathcal{C}^{\pm}_{k,\lambda}$. Moreover, we explicitly construct a circle in terms of elliptic theta functions. This result implies the existence of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial ones has remained an open problem. Since an element of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a Kaleidocycle with $k$-tetrahedra in a specific position (Section 2.1) and its motion is described by a path in the configuration space, our main result provides an explicit construction and its periodic motion for Kaleidocycles of any $k\geq 6$.
@@
The key observation is to identify a Kaleidocycle with a closed spatial polygonal curve formed by connecting the midpoints of the hinge edges. Then, the angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the signed curvature angle (§2.1). An element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a sequence of binormals. In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle $\lambda$. In our construction, it plays the central role to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
@@
It was known in [21] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (see §3.2). The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12]5). Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 18].
@@
Both views combined, a flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ is explicitly integrated using elliptic theta functions. Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, the Kaleidocycle.
@@
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. They are identified with a certain class of discrete spatial curves. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. In Section 5, a family of time-parametrised curves is constructed by introducing an additional parameter $t$. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the curve and show connections to previous research. Section 6 investigates a necessary and sufficient condition for the resulting curve to be closed for every $t$. We prove there exists a set of parameters that satisfies the closure condition for a given $k\geq 6$, which proves the existence of Kaleidocycles. Section 7 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present a conjecture (Conjecture 7.1) suggested by numerical experiments: the Kaleidocycles constructed in this paper seem to be so-called Möbius Kaleidocycles.
@@ 2. Algebraic and geometric models of Kaleidocycle
2.1. Configuration space of a Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein). However, there is no literature that discusses the existence of such objects for arbitrary $k$.
We first define a generalisation of the Kaleidocycle following [20]. We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring. Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces. Let $l_{01},l_{12},l_{02}$ be the lengths of the sides of the triangle face, with the edge of length $l_{01}$ serving as a hinge (see Figure 1 Right). Now, a natural way to model a Kaleidocycle is as a realisation of an edge-weighted graph. Let $(V,E)$ be the graph formed by the vertices and edges of the tetrahedra in a Kaleidocycle and $\ell:E\to\mathbb{R}$ be the edge length. A realisation of $(V,E,\ell)$ in $\mathbb{R}^{3}$ is a map $\phi:V\to\mathbb{R}^{3}$ satisfying $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$.
In general, the collection of all possible realisations of an edge-weighted finite graph $(V,E,\ell)$ in $\mathbb{R}^{d}$ forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$. This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [11, 23, 27].
More concretely, a state of a $k$-Kaleidocycle of type $(l_{01},l_{12},l_{02})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k+1})\in(\mathbb{R}^{3})^{2k+2}$ of the quadratic equations
@@ 2.1. Configuration space of a Kaleidocycle
$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12},$
for $0\leq n\leq k-2$ and $(x_{2k},x_{2k+1})=(x_{0},x_{1})$ or $(x_{1},x_{0})$. To remove the translational freedom, we demand $x_{0}+x_{1}=0$.
Notice that we have two combinatorially distinct cases: length $l_{02}$ edges form a $k$-cycle ($x_{2k}=x_{0}$) or $2k$-cycle, going around the ring twice ($x_{2k}=x_{1}$).
Remark 2.1.
@@
The solution space of (2.1) admits an action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $3\cdot 2k-5k-3-\dim(O(3))=k-6$, where $5k$ is the number of edges and $3$ comes from the equation $x_{0}+x_{1}=0$. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [9]. However, this estimate is sometimes violated in interesting examples. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 0, giving an example of an overconstrained system  [9]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [22]. [^1] Some of the findings of [22] are also published in [31]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of an underconstrained linkage mechanism whose configuration space has a smaller dimension than the expected one.
We use the following elementary geometric property of equifacial tetrahedron:
Lemma 2.2.
An equifacial tetrahedron whose faces are congruent triangles with edge-lengths $(l_{01},l_{12},l_{02})$ may be realised within a rectangular cuboid of side lengths
$\sqrt{\frac{l_{01}^{2}+l_{12}^{2}-l_{02}^{2}}{2}},\quad\sqrt{\frac{l_{12}^{2}+l_{02}^{2}-l_{01}^{2}}{2}},\quad\sqrt{\frac{l_{02}^{2}+l_{01}^{2}-l_{12}^{2}}{2}}$
so that tetrahedron edges are diagonals of the faces. Moreover, the twist angle between the pair of opposite edges of length $l_{01}$ is given by the relation
$\cos(\lambda)=\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}.$
We will see the solution space of Equation 2.1 is identified with $\mathcal{C}^{\pm}_{k,\lambda}$.
Proposition 2.3.
The space of real solutions to Equation 2.1 is $O(3)$-equivariantly homeomorphic to $\mathcal{C}^{\pm}_{k,\lambda}$ with $\lambda=\arccos\left(\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}\right)$.
Proof.
Given $(x_{0},\ldots,x_{2k+1})$ satisfying Equation 2.1, we define $b_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$. By 2.2, the inner product $\langle b_{n-1},b_{n}\rangle$ is constant $\cos\lambda$ for $1\leq n<k$.
Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge. There exists $l>0$ such that $l|b_{n+1}\times b_{n}|=|\gamma_{n+1}-\gamma_{n}|$. By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$. Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=l\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron. We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$. Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
It is clear that the above constructions are mutually inverse. ∎
While the shape of the tetrahedron is determined by the triangle edge lengths $(l_{01},l_{12},l_{02})$, the configuration space depends only on the hinge angle $\lambda$. We call $\lambda$ the torsion angle.
We call the (possible non-flat) polygon, or closed discrete spatial curve, defined by $\gamma_{n}$ the central curve of the Kaleidocycle.
Depending on the sign of $b_{k}=\pm b_{0}$, we say the Kaleidocycle is oriented (resp. non-oriented).
@@
Figure 1. (Left) The classical 6-Kaleidocycle. (Centre) Kaleidocycle as a discrete curve with its framing (Right) Net of a Kaleidocycle with labelled vertices and edge lengths.
@@ 2.2. Discrete spatial curve of a constant torsion angle
We saw that a discrete spatial curve $\gamma_{n}$ is associated with a Kaleidocycle. We review some basics on discrete curves.
A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals. Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\left|\widetilde{B}_{n}\times\widetilde{B}_{n-1}\right|\neq 0$ for all $n\in\mathbb{Z}$. A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows. Define the tangent $T:\mathbb{Z}\rightarrow S^{2}$ and the normal $\widetilde{N}:\mathbb{Z}\rightarrow S^{2}$ by
@@
(2.4) $\begin{split}\langle T_{n},T_{n-1}\rangle=\cos\kappa_{n},\quad\langle\widetilde{N}_{n},T_{n-1}\rangle&=-\sin\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
@@
Note that our Frenet frame coincides with the usual one only up to sign. We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes through the state where the curvature vanishes.
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen.
In this language, a Kaleidocycle is equivalent to a discrete closed ($\gamma_{k}=\gamma_{0})$ spatial curve with constant torsion angle $\lambda$ determined by the binormal $\widetilde{B}_{n}=b_{n}$. Our approach is to utilise this identification.
The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By the Călugăreanu-Fuller-White theorem [1], the writhing number of the curve must remain constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.